Reports 1-1 of 1 Clear search Modify search
AOS (Baffles & Beam dumps)
simon.zeidler - 21:59 Wednesday 14 November 2018 (6985) Print this report
NAB IYA finalization part 1

Akutsu, Yano, K. Tanaka, Simon

We started the final alignment check after the NAB has been inserted into the chamber. However, there were several things that needed to be done in advance:

  • we actually devided us into two teams in the morning
    • one team (Yano-san and I) were attaching the magnet-dampers to the NAB
    • the other team (Akutsu-san and Tanaka-kun) was bringing the OpLev-setup from IXA to IYA (for taking measurements)
  • Before we started with the dampers, we set all the missing M5 screws to fix the NAB onto the tower. As reported last week (6916), we could set those screws only on one side of the NAB
  • Yano-san and I recognized that the screws from several dampers became a little bit loosen. So, we tightened all screws again and attached the damper to the NAB.
  • we found that the gap between the centers of the damper-magnet holder's spring-blades (where some flat-head screws are existing) is extremely small (<0.5mm, see picture 2). After some email traffic with Ikenoue-san, we learned that we would have to insert spacers in between the outer clamps and the spring-clamps to make the gap wider. Fortunately, the spacers are identical with the upper damper-spring clamp and we still had the two older damper-magnet holders in IXA having 8 of those clamps in total (a perfect match with our needs). So team #2 brought us the old holders and we took the clamps and made them to spacers for the actual magnet-dampers. The result is quite good: the gap is now ~1mm! (see picture 3)
  • while team #2 brought the OpLev parts (pylon, breadboard with light-source, etc.), Yano-san and I attched a SUS plate around the center-position of the NAB on the -X side, where the OpLev is supposed to be set

In the afternoon, we released the baffle and checked the alignment. We found a strong pitch of the baffle (picture 4 shows the effect on a bubble level), which is very likely triggered by the PD-holders that have been attached to the baffle last week.
In order to bring the baffle back to a kind of balanced state, we had to remove 400g of the attached balance masses (on the same side as the masses, the PD-holders are attached; the result can be seen in picture 5).
A futher check of the position of the baffle with respect to the main beam-line (or better: to some indications inside and outside the chamber) has been done after the baffle-rebalancing:

  • the baffle itself is quite centered vertically; the relative divergence to the beam-line height is ~1mm, or less
  • the horizontally alignment with respect to the indication lines inside the chamber is less centered; it is hard to put in numbers but there is s slight shift of the baffle toward +X by ~2mm

Akutsu-san also placed a small viewport window for the OpLev, while Tanaka-kun prepared cabels and electronics for it.
We also prepared already the laser-displacement sensors for the measurement of the eigenmodes and the vibration-isolation, which will taken place tomorrow.

Images attached to this report
Comments to this report:
tomotada.akutsu - 23:59 Wednesday 14 November 2018 (6987) Print this report

ICF70 oplev viewport window S/N002 has been installed at IYA.

Search Help